# The nature of ASASSN-24fw’s occultation: modelling the event as dimming by optically thick rings around a sub-stellar companion

### v1

@@ 2.7 Spectral Energy Distribution
We form the spectral energy distribution of ASASN-24fw using publicly available pre-dimming photometry from various surveys. For this, we obtained single epoch photometric data in the optical from SDSS (Almeida et al., 2023), PANSTARRS DR2 (Flewelling, 2018), Gaia DR3 (Gaia Collaboration et al., 2023), HST (Whitmore et al., 2016), and in the infrared 2MASS (Skrutskie et al., 2006) and WISE (Wright et al., 2010). We corrected these magnitudes (as well as our observed spectra during dimming) for extinction using the mean value of pre-dimming reddening from Schlafly & Finkbeiner (2011) reddening maps. These maps are updated from Schlegel et al. (1998) reddening maps and use the standard Fitzpatrick (1999) reddening law. We find a relatively low value of E(B-V) 0.05 for ASASSN-24fw and a clear presence of infrared excess in the WISE points as shown in the first panel of Figure (2). In the second panel of the Figure (2), we can also see the SED formed using the spectrum observed during the dimming. In the following sections, we describe in detail about our SED fitting methods.
2.8 Extinction calculations using LCOGT data
Our aim of obtaining multi-band LCOGT observations was to monitor the variation in color of ASASSN-24fw during dimming. The variation in color can then be used to compute the specific reddening vector and hence the composition of the occulting material. Unfortunately, our post-dimming baseline of ASASSN-24fw using LCOGT observations in the four SDSS filters is not very long. However, it is still enough to compute the reddening vector and hence constrain the primary composition of the occulting material. We generate color-color plots using g’, r’, i’, and z’s) bands and use optool (Dominik et al., 2021) to calculate the opacities (= +) values for a range of dust grain compositions and two size distributions. We used the built-in refractive indices from optool to generate the values for materials including astronomical silicate (Draine, 2003), various crystalline and amorphous iron-magnesium silicates (Dorschner et al., 1995; Jaeger et al., 1998; Fabian et al., 2001; Suto et al., 2006), carbon-bearing species (Draine & Malhotra, 1993; Henning & Stognienko, 1996), quartz (Kitamura et al., 2007), and iron particles (Henning & Stognienko, 1996).
We consider a log-normal grain distribution with a width of 0.5 dex and a free peak size, considering contributions from sizes between 0.1 and 1.1m in steps of 0.01m when calculating , determined from the complex refractive indices assuming Mie theory (Mie, 1908). For a Sun-like star, the typical blowout size for a dust grain is 0.5m (Krivov, 2010), so we include transient dust with this range of grain sizes. The optical depth and extinction of a dust cloud that fully covers the stellar disc are then calculated from the values for each combination of composition and size distribution, which are then compared to the measured reddening vectors by least-squares fit to identify those which best match the observations. For ASASSN-24fw, we find that no combination of compositions available in optool and grain size produces a reddening vector that fits the data. This indicates that the occulting material comprises of large bodies and may also contain gas.
@@ 3 Nature of ASASSN-24fw
| Teff | |
| log(g) | |
The Gaia DR3 (Gaia Collaboration et al., 2023) properties of ASASN-24fw are summarized in Table (1). Here we also show the PANSTARRS, 2MASS, and ALLWISE IDs. The G-band absolute magnitude was determined from the extinction () corrected apparent G magnitude by subtracting the distance modulus, which was calculated from the distance of 104417 pc derived from the trigonometric parallax. Here, we have also included the 3D peculiar velocity computed using the equations by Reid et al. (2009), and Poleski (2013). We find that ASASSN-24fw belongs to the Milky Way’s thick disc distribution. We use galpy (bovy2015), a Python package to compute the orbit of ASASSN-24fw in Milky Way in the past 50 Ma. galpy supports orbit integration in a variety of gravitational potentials, evaluating and sampling various distribution functions, and the calculation of action-angle coordinates for all static potentials. We find that the orbit of ASASN24fw does not intersect with any star formation region. Therefore this tells us that although ASASN-24fw has a persistent infrared excess, it is unlikely to be a Young Stellar Object (YSO) and is more likely to be on main-seqenuce.
@@ 3.1 Search for intrinsic variability
We searched for variability related to inherent stellar phenomena like star-spots, accretion, or rotation which is high in pre-main-sequnece (PMS) stars in the TESS lightcurve of ASASSN-24fw. We find a peak at 0.41 d with a very small false-alarm probability. But the phase-folded lightcurve about this period is flat indicating that the frequency may not have any real significance. Since we do not have TESS lightcurve during dimming episode, we could not check the during dimming periodogram. But, instead we used ASAS-SN lightcurve to find repeating signals in the lightcurve of both the pre- and during dimming phase. Here we saw a peak at 1 day in both the periodograms which is actually common in ground-based photometry, and does not reflect any variability related to the star. As YSOs have high magnetic field, accretion and star spot activity, our analysis re-iterates that ASASSN-24fw is unlikely to be on pre-main sequence and is in fact more evolved.
3.2 Fundamental stellar parameters using Stellar Evolutionary Models applied pre-dimming SED
In the pre-dimming SED, we clearly see an infrared excess particularly in the mid-infrared WISE data points (the first panel of Figure (2)). We fit this SED with Castelli-Kurucz models222We have used VOSA (VOSA v7.5, bayo2008) https://svo2.cab.inta-csic.es/theory/vosa/ to perform this analysis (Castelli & Kurucz, 2003) and find the best fitting model has [Teff,log(g),[Fe/H]] = [6250,5.0,-0.5]. This model is shown in cyan color in all the panels of Figure (2).
We then applied the Bayesian inference code of del Burgo & Allende Prieto (2016, 2018) on a grid of stellar evolution models constructed from the PARSEC v1.2S tracks (Bressan et al., 2012; Chen et al., 2014, 2015; Tang et al., 2014) to derive the fundamental stellar parameters of ASASSN-24fw. This choice is supported by the good statistical match obtained for detached eclipsing binaries, particularly for stars on the main-sequence, where the measured dynamical masses and corresponding predictions are consistent on average to within 4% (del Burgo & Allende Prieto, 2018).
The grid of PARSEC v1.2S models arranged in this code consisted of ages ranging from 2 to 13,800 Ma in steps of 5%, and [M/H] from -2.18 to 0.51 in steps of 0.02 dex. We adopted the photometric passband calibrations of Riello et al. (2021), and the zero points of the VEGAMAG system. We fed our code by the following three input parameters: the absolute magnitude , the color (also adjusted for reddening and extinction), and the iron to hydrogen abundance (assumed to be solar [Fe/H]= dex.
@@ 3.2 Fundamental stellar parameters using Stellar Evolutionary Models applied pre-dimming SED
In Figure (2), we show the two different Kurucz models corresponding to the stellar atmospheric parameters that we just obtained. In addition to the cyan colored model that we discussed earlier, the green colored Kurucz model corresponds to (Teff, log(g), [Fe/H]) = (6000, 4.0, -0.5) and the red colored Kurucz model corresponds to (Teff, log(g), [Fe/H]) = (6500, 5.0, 0.0) i.e. the two models that we obtained by using the code of del Burgo & Allende Prieto (2016, 2018).
In the second panel, we show the SED of the star during dimming formed using our observed spectra. We see that all the three Kurucz models fit well to the continuum but the main difference comes while checking the fit to the peak of the SED especially Ca II HK absorption lines. In the same plot, we have also shown an M8 brown dwarf NIR spectrum template overplotted on top of the observed Palomar NIR spectrum. This NIR template spectrum is taken from the NIRSPEC Spectroscopic Brown Dwarf Survey (McLean et al., 2007) which has a collection of spectral templates for the M, L, and T type brown dwarfs. We can clearly see that the template matches the continuum and there is an excess infrared flux from the end of H-band. This hints towards the elevated levels of infrared excess during dimming perhaps caused by circum-planetary rings to the brown dwarf. Since we do not have any mid-infrared observations during dimming, we cannot compare the pre-dimming and post-dimming fluxes for mid-infrared wavelengths. Finally in the bottom-panel we show the residual flux after subtracting the three model spectra from the observed spectra.
3.3 Blackbody fit
@@
The emission in the mid-infrared and longer wavelengths comes from the circum-stellar dust. This dust particles absorb the incident stellar flux and re-emit into the longer wavelength regions. Thus the nature of the excess flux can tell us about the nature of the circum-stellar material. We can assume that the dust particles act as perfect black bodies and model the deviant points corresponding to the excess flux in the SED using a solid blackbody model. In Figure (2), it is visible that the excess flux actually manifests from the 2MASS band. So instead of a single component, we fit a two-component blackbody model. The warm component of our best fitted models gives us an effective temperature of 1072K for the primary and 387K for the secondary colder component. An effective temperature of 1072K indicates a warm circum-stellar material within 0.2 au radius from the star which is well below the dust sublimating temperature while an effective temperature of 387K hints at a colder disc located beyond 1.0 au from the star may consist of larger and rocky bodies. The two component black-body model is shown as a black dashed line passing through the WISE points.
@@ 3.4 Spectral Analysis
The spectra that we have obtained were during the dimming episode of ASASN-24fw. Thus they are a mixture of the stellar spectra and the occulter absorption sprectra. It is therefore possible that, along with the continuum, the occulter may have affected the lines in the stellar spectra. For example, the variation in the Hydrogen series lines and the Ca II HK lines in Figure (3) in the LRIS spectrum taken on 31 October 2025 and HiRES spectrum taken on 28 January 2025 indicates that the gaseous component in the occulter is affecting the stellar spectra. Further, in the HiRES spectrum, we see that the age indicator Li line is not detected (or very weak at the level of noise), Fe, and Mg lines are moderately detected. These features as well as the shape and depth of the Ca II HK lines are normally expected for an F-type main-sequence star. Unlike YSOs, we do not see any emission line in our spectra. However, we see a "reversal" feature or shape in the H absorption line of the HiRES spectrum. Whether this is astrophysical or due to spectral noise is not very clear from the single spectrum. However, to find the strength of this line, we calculated the equivalent width of the absorption () and central emission () component of the H line. For this, we fitted a Lorentzian profile to the absorption line by masking the emission line corrected for barycentric motion. We find that 93Å and 7.8Å and the radial velocity of the emission line is -7.14 km/s.
@@ 4.1 Lightcurve Modelling
While we focused on the earlier on characterizing the star, now we focus on understanding the occulter. The shape of the lightcurve can tell us a lot about its nature eg. small and regular dips can be attributed to exoplanets, while the irregular episodes of dimming can be attributed to patchy or fragmented circum-stellar material. However, a long period and large scale dimming requires an object comparable to the size of the star. A single dark body comparable to the size of the star is physically impossible. But, the occulting body could be considered as a system of central body with massive circum-planetary rings and depending on the orientation of the ring system, a large portion of the star light can be blocked for longer duration. We proceed with lightcurve modelling in two steps; first we model the lightcurve using a parametric template that matches the shape of the lightcurve. This helps us to derive some basic parameters like the center of dimming, characteristic timescale, and lightcurve depth. In the next step, we can find a ring system geometry which can best explain the shape of the lightcurve. Here the information from the first model helps in setting the priors for Bayesian analysis used to maximize the likelihood of the fit. This is explained in the following sections.
@@ 4.1.1 Transit-Profile model
where () is the model brightness, is the maximum depth of the dimming, is the time of center of dimming, is the duration of the dimming, and is a parameter that defines the shape of the dimming. In case of ASASSN-24fw, should be high because of the flatness of the occulted lightcurve. ASASN-24fw dropped to 16.75 magnitudes during dimming which is close to the detection limit of 17.5 mag of ASAS-SN camera. Therefore, the lightcurve would be affected by noisy data points. In order to model the long-term trends in the lightcurve and to handle the noisy data points, in the next step we used the non-parametric Gaussian Process regression model. This was implemented through a Python package called george (Ambikasaran et al., 2015), with a squared exponential kernel (ExpSquaredKernel) given by:
@@
where and are the kernel hyper-parameters. While the former hyper-parameter is the amplitude of the correlated noise in the whole lightcurve, the latter is the characteristic timescale of the noise in the whole lightcurve. Using these two hyper-parameters, george will try to find a mean model through the distributed data points. In this process, it can highlight any trends in the data which can otherwise be missed. Therefore, we combine these two models to capture the dimming and non-periodic trends in the lightcurve. We call our combined model as Transit-Profile model. We also performed the Bayesian analysis to maximize the likelihood of the fit and calculated the posterior distribution of our parameters. For this, we used the emcee (Foreman-Mackey et al., 2013) Python package which implements the Affine Invariant Markov Chain Monte Carlo (MCMC) Ensemble sampler algorithm introduced by Goodman & Weare (2010).
@@
Our best-fitted Transit-Profile model is shown in the first panel of Figure (4) and the parameters with their 16th, 50th, and 84th percentile values listed in Table (4). This fit tells use few things; first, the star started dimming gradually 150 days (roughly in March 2024) much before the actual alert was issued during the ingress. The second phase of dimming started around 02 September 2024 continued up to 13.66 magnitudes around 16 September 2024. Finally, the third phase of ingress caused a drastic and sudden drop in magnitude and the brightness reduced to 16.70 around 02 October 2024. This indicates that there are multiple layers of varying densities and optical thickness of material around the occulting body, possibly hinting towards a gas-giant body with massive rings.
@@
To check if our Transit-Profile model is really capturing various trends in the data or is just over-fitting the lightcurve. For this, we repeat the transit-profile model fitting on the ATLAS survey lightcurve. The depth of ATLAS camera is higher than the ASAS-SN camera and hence the data would be less noisy. We see that the wiggles or the "micro-trends" in the fit have reduced but the three phases or the "macro-trends" of ingress are similarly fitted (see the red and green colored data points and the model passing through them in the Figure (4)). This indicates that the long-term or the "macro-trends" in the lightcurve are really present and discovered using our method.

### v2

@@ 2.7 Spectral Energy Distribution
We form the spectral energy distribution of ASASSN-24fw using publicly available pre-dimming photometry from various surveys. For this, we obtained single epoch photometric data in the optical from SDSS (Almeida et al., 2023), PANSTARRS DR2 (Flewelling, 2018), Gaia DR3 (Gaia Collaboration et al., 2023), HST (Whitmore et al., 2016), and in the infrared 2MASS (Skrutskie et al., 2006) and WISE (Wright et al., 2010). We corrected these magnitudes (as well as our observed spectra during dimming) for extinction using the mean value of pre-dimming reddening from Schlafly and Finkbeiner (2011) reddening maps. These maps are updated from Schlegel et al. (1998) reddening maps and use the standard Fitzpatrick (1999) reddening law. We find a relatively low value of E(B-V) 0.05 for ASASSN-24fw and a clear presence of infrared excess (around 10%) in the WISE points as shown in the first panel of Figure (2). In the second panel of the Figure (2), we can also see the SED formed using the spectrum observed during the dimming. In the following sections, we describe in detail about our SED fitting methods.
@@ 3 Nature of ASASSN-24fw
| eff | |
| Metallicity ([M/H]) |
The Gaia DR3 (Gaia Collaboration et al., 2023) properties of ASASSN-24fw are summarized in Table (1). The G-band absolute magnitude was determined from the extinction () corrected apparent G magnitude by subtracting the distance modulus, which was calculated from the distance of 104417 pc derived from the trigonometric parallax. Here, we have also included the 3D peculiar velocity computed using the equations by Reid et al. (2009), and Poleski (2013). We find that ASASSN-24fw belongs to the Milky Way’s thick disc distribution. We use galpy (Bovy, 2015), a Python package to trace the orbit of ASASSN-24fw in Milky Way in the past 50 Ma. galpy supports orbit integration in a variety of gravitational potentials, evaluating and sampling various distribution functions, and the calculation of action-angle coordinates for all static potentials. We find that the orbit of ASASSN-24fw does not intersect with any star formation region. Therefore this tells us that although ASASSN-24fw has a persistent infrared excess, it is unlikely to be a Young Stellar Object (YSO) and is more likely to be on main-sequence.
@@ 3.1 Search for intrinsic variability
We searched for variability related to inherent stellar phenomena like star-spots, accretion, or rotation which is high in pre-main-sequence (PMS) stars in the TESS lightcurve of ASASSN-24fw. We find a peak at 0.41 d with a very small false-alarm probability. But the phase-folded lightcurve about this period is flat indicating that the frequency may not have any real significance. Since we do not have TESS lightcurve during dimming episode, we could not check the during dimming periodogram. But, instead we used ASAS-SN lightcurve to find repeating signals in the lightcurve of both the pre- and during dimming phase. Here we saw a peak at 1 day in both the periodograms which is actually common in ground-based photometry, and does not reflect any variability related to the star. As YSOs have high magnetic field, accretion and star spot activity, our analysis suggests that ASASSN-24fw is unlikely to be on pre-main-sequence and could be more evolved.
3.2 Fundamental stellar parameters using Stellar Evolutionary Models applied to pre-dimming SED
In the pre-dimming SED, we clearly see an infrared excess particularly in the mid-infrared WISE data points (the first panel of Figure (2)). We fit this SED with Castelli-Kurucz models333We have used VOSA (VOSA v7.5, Bayo et al., 2008) https://svo2.cab.inta-csic.es/theory/vosa/ to perform this analysis (Castelli and Kurucz, 2003) and find the best fitting model has [Teff,log(g),[Fe/H]] = [6250,5.0,-0.5]. This model is shown in cyan color in all the panels of Figure (2).
We then applied the Bayesian inference code of del Burgo and Allende Prieto (2016, 2018) on a grid of stellar evolution models constructed from the PARSEC v1.2S tracks (Bressan et al., 2012; Chen et al., 2014, 2015; Tang et al., 2014) to derive the fundamental stellar parameters of ASASSN-24fw. This choice is supported by the good statistical match obtained for detached eclipsing binaries, particularly for stars on the main-sequence, where the measured dynamical masses and corresponding predictions are consistent on average to within 4% (del Burgo and Allende Prieto, 2018).
The grid of PARSEC v1.2S models arranged in this code consisted of ages ranging from 2 to 13,800 Ma in steps of 5%, and [M/H] from -2.18 to 0.51 in steps of 0.02 dex. We adopted the photometric passband calibrations of Riello et al. (2021), and the zero points of the VEGAMAG system. We fed our code by the following three input parameters: the absolute magnitude , the color (also adjusted for reddening and extinction using the values provided by Gaia Data Archive (Gaia Collaboration et al., 2023)), and the iron to hydrogen abundance (assumed to be solar [Fe/H]= dex).
@@ 3.2 Fundamental stellar parameters using Stellar Evolutionary Models applied to pre-dimming SED
In Figure (2), we show the two different Kurucz models corresponding to the stellar atmospheric parameters that we just obtained. In addition to the cyan colored model that we discussed earlier, the green colored Kurucz model corresponds to [Teff, log(g), [Fe/H]] = [6000, 4.0, -0.5] and the red colored Kurucz model corresponds to [Teff, log(g), [Fe/H]] = [6500, 5.0, 0.0] i.e. the two models that we obtained by using the code of del Burgo and Allende Prieto (2016, 2018).
3.3 Blackbody fit to the pre-dimming SED
@@
The emission in the mid-infrared and longer wavelengths comes from the circumstellar dust. This dust particles absorb the incident stellar flux and re-emit into the longer wavelength regions. Thus the nature of the excess flux can tell us about the nature of the circumstellar material. We can assume that the dust particles act as perfect black bodies and model the deviant points corresponding to the excess flux in the SED using a solid blackbody model. In Figure (2), it is visible that the excess flux actually manifests from the 2MASS band. So instead of a single component, we fit a two-component blackbody model. The warm component of our best fitted models gives us an effective temperature of 1100K for the primary and 400K for the secondary colder component. An effective temperature of 1100K, which is well below dust sublimation temperature, indicates a warm circumstellar material within 0.2 au radius from the star which is well below the dust sublimating temperature while an effective temperature of 387K hints at a colder disc located beyond 1.0 au from the star may consist of larger and rocky bodies. The two component black-body model is shown as a black dashed line passing through the WISE points.
3.4 SED using the spectral observations during dimming
In the second panel of Figure (2), we show the SED of the star during dimming formed using our observed spectra. We see that all the three Kurucz models fit well to the continuum but the main difference comes while checking the fit to the peak of the SED especially Ca II H&K absorption lines which are deepest among all the lines. In the same plot, we have also shown an M8 brown dwarf NIR spectrum template overplotted on top of the observed Palomar NIR spectrum and the residual flux after subtracting the three model spectra from the observed spectra. This NIR template spectrum is taken from the NIRSPEC Spectroscopic Brown Dwarf Survey (McLean et al., 2007) which has a collection of spectral templates for the M, L, and T type brown dwarfs. We can clearly see that the template matches the initial part of the near-infrared continuum and there is an excess infrared flux from the end of H- band. This hints towards the elevated levels of infrared excess during dimming perhaps caused by circum-planetary rings to the brown dwarf. We note that this excess is in the near-infrared wavelengths unlike the WISE wavelengths which span the mid-infrared region. Thus, the excess that we observe during dimming is most likely to be related to the occulter rather than the blackbodies that we fitted to the WISE points. Since we do not have any mid-infrared observations during dimming, we cannot compare the pre-dimming and post-dimming fluxes for mid-infrared wavelengths.
@@ 3.5 Spectral Analysis
The spectra that we have obtained were during the dimming episode of ASASSN-24fw. Thus they should be a mixture of the stellar spectra and the occulter spectra. It is therefore possible that, along with the continuum, the occulter may have affected the lines in the stellar spectra. For example, the variation in the hydrogen series lines and the Ca II HK lines in Figure (3) in the LRIS spectrum taken on 31 October 2025 and HiRES spectrum taken on 28 January 2025 indicates that the occulter may be affecting the stellar spectra. Further, in the HiRES spectrum, we see that the age indicator Li line is not detected (or very weak at the level of noise), Fe, and Mg lines are moderately detected. These features as well as the shape and depth of the Ca II HK lines are normally expected for an F-type main-sequence star. Unlike YSOs, we do not see any emission line in our spectra. However, we see a central "reversal" feature or shape in the H absorption line of the HiRES spectrum. Whether this is astrophysical or due to spectral noise is not very clear from the single spectrum. However, to find the strength of this line, we calculated the equivalent width of the absorption () and central emission () component of the H line. For this, we fitted a Lorentzian profile to the absorption line by masking the emission line corrected for barycentric motion. We find that 93Å and 7.8Å and the radial velocity of the emission line in the heliocentric frame is -7.14 km/s. In contrast, Zakamska et al. (2025) find that a two gaussian fit to their H alpha emission yields an EW 1.98 Å and the velocity centroid of the emission line is -19 km/s in the heliocentric frame. This value is different from ours because the epoch of spectra as well as instruments are different.
@@ 4.1 Lightcurve Modelling
While we initially focused on characterizing the star, we now focus on understanding the occulter. The shape of the lightcurve can tell us a lot about its nature eg. small and regular dips can be attributed to exoplanets, while the irregular episodes of dimming can be attributed to patchy or fragmented circumstellar material. However, a long period and large scale dimming requires an object comparable to the size of the star. A single dark body comparable to the size of the star is physically impossible. But, the occulting body could be considered as a system of central body with massive circumplanetary rings and depending on the orientation of the ring system, a large portion of the star light can be blocked for longer duration. We proceed with lightcurve modelling in two steps; first we model the lightcurve using a parametric template that matches the shape of the lightcurve. This helps us to derive some basic parameters like the center of dimming, characteristic timescale, and lightcurve depth. In the next step, we can find a ring system geometry which can best explain the shape of the lightcurve. Here the information from the first model helps in setting the priors for Bayesian analysis used to maximize the likelihood of the fit. This is explained in the following sections.
@@ 4.1.1 Transit-Profile model
where () is the model brightness, is the maximum depth of the dimming, is the time of center of dimming, is the duration of the dimming, and is a parameter that defines the shape of the dimming. In case of ASASSN-24fw, should be high because of the flatness of the occulted lightcurve. ASASSN-24fw dropped to 16.75 magnitudes during dimming which is close to the detection limit of 17.50 mag of ASAS-SN camera. Therefore, the lightcurve would be affected by noisy data points. In order to model the long-term trends in the lightcurve and to handle the noisy data points, in the next step we used the non-parametric Gaussian Process regression model. This was implemented through a Python package called george (Ambikasaran et al., 2015), with a squared exponential kernel (ExpSquaredKernel) given by:
@@
where and are the kernel hyper-parameters. While the former hyper-parameter is the amplitude of the correlated noise in the whole lightcurve, the latter is the characteristic timescale of the noise in the whole lightcurve. Using these two hyper-parameters, george will try to find a mean model through the distributed data points. In this process, it can highlight any trends in the data which can otherwise be missed. Therefore, we combine these two models to capture the dimming and non-periodic trends in the lightcurve. We call our combined model as Transit-Profile model. We also performed the Bayesian analysis to maximize the likelihood of the fit and calculated the posterior distribution of our parameters. For this, we used the emcee (Foreman-Mackey et al., 2013) Python package which implements the Affine Invariant Markov Chain Monte Carlo (MCMC) Ensemble sampler algorithm introduced by Goodman and Weare (2010).
@@
Our best-fitted Transit-Profile model is shown in the first panel of Figure (4) and the parameters with their 16th, 50th, and 84th percentile values listed in Table (4). This fit tells use few things; first, the star started dimming gradually 150 days (roughly in March 2024) much before the actual alert was issued during the ingress. The second phase of dimming started around 02 September 2024 continued up to 13.66 magnitudes around 16 September 2024. Finally, the third phase of ingress caused a drastic and sudden drop in magnitude and the brightness reduced to 16.70 around 02 October 2024. This indicates that there are multiple layers of varying optical depth of material around the occulting body, possibly hinting towards a gas-giant body with optically thick rings. To check whether the gradual decrease in brightness is due to systematics, we downloaded the ASAS-SN lightcurves for all the stars within 3 radius of ASASSN-24fw that had magnitudes 11 < < 15. We found four such stars. But they were invariant at 0.10 mag level (3 level). This rules out the fact that the Phase I of the ingress is due to telescope systematics but is due to the outer layers of the occulter system.
@@
To check if our Transit-Profile model is really capturing various trends in the data or is just over-fitting the lightcurve. For this, we repeat the transit-profile model fitting on the ATLAS survey lightcurve. The photometric limit of ATLAS camera is deeper than the ASAS-SN camera and hence the data would be less noisy. We see that the wiggles or the "micro-trends" in the fit have reduced but the three phases or the "macro-trends" of ingress are similarly fitted (see the red and green colored data points and the model passing through them in the Figure (4)). This indicates that the long-term or the "macro-trends" in the lightcurve are really present and discovered using our method.